The length of the rectangular plot is 3 times the breadth of the plot is b.Form the equation for its length and perimeter?

Answer by ikleyn(52778)

.
The length of the rectangular plot is 3 times the breadth of the plot is b.Form the equation for its length and perimeter?
~~~~~~~~~~~~~~~~~~~~~~~

The length L = 3b

Perimeter = L + b + L + b = 3b + b + 3b + b = 8b.